I'm trying to help a friend out with his dirty harry. It's exhibiting some switch behaviors I haven't seen before, when you think it's all about the optics, then one of the physical ones throws you for a loop.
When everything is connected, the switch test shows rows 1,3,6 and 7 all closed (pic1). If we disconnect the optics by pulling off J3 from the 16 Opto Switch board A-169998, all of those rows return to normal (except the opticals that show show normally closed of course). We've traded out that board from my ST:TNG and the behaviors don't change, so we don't believe the problem is in the board.
When we pull the J3 from that board, and then test individual switches, switches 11 through 28 function fine (columns 1 and 2), when we go to test any other physical switches, for example - #61 Left sling, it will show as row 1 completely closed. (pic 2)
I *think* the switch matrix ties back to U18,U19 and U20 on the Security CPU board (Pic3, lower left) - U20 was recently replaced, should we replace U18 and U19, LM339's?
Thanks for the troubleshooting directions!